OpenStudy (anonymous):

my answer is b Simplify: (3x+2y)(3x-2y)= 9x2-4y2 9x2+4y2 9x2+12xy+4y2 None of the above

jimthompson5910 (jim_thompson5910):

b is incorrect

OpenStudy (anonymous):

a is correct

jimthompson5910 (jim_thompson5910):

similar example (6x-7y)(6x+7y) = (6x)^2 - (7y)^2 (6x-7y)(6x+7y) = 36x^2 - 49y^2

OpenStudy (anonymous):

This is a situation involving Difference of squares:\[(a+b)(a-b)=a^2-b^2 \rightarrow (3x+2y)(3x-2y)=..?\] @InsanelyChaotic

OpenStudy (anonymous):

B!

jimthompson5910 (jim_thompson5910):

B) 9x^2+4y^2 is incorrect

OpenStudy (anonymous):

Let me check..9x2-4y2 <--- answer

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

thank you jimmy neutron

jimthompson5910 (jim_thompson5910):

lol yw

OpenStudy (anonymous):

bye jimmy

jimthompson5910 (jim_thompson5910):

cya